# The Shape of Word Embeddings: Recognizing Language Phylogenies through Topological Data Analysis

### v1

@@ Abstract
Word embeddings represent language vocabularies as clouds of $d$-dimensional points. We investigate how information is conveyed by the general shape of these clouds, outside of representing the semantic meaning of each token. Specifically, we use the notion of persistent homology from topological data analysis (TDA) to measure the distances between language pairs from the shape of their unlabeled embeddings. We use these distance matrices to construct language phylogenetic trees over 81 Indo-European languages. Careful evaluation shows that our reconstructed trees exhibit strong similarities to the reference tree.
@@ 1 Introduction
Word embeddings are well-established objects of interest in natural language processing, being $d$-dimensional vector representations that capture the semantics of each vocabulary word. The vocabulary of language $L$ can thus be viewed as a cloud of points, whose geometric and structural properties encode considerable information about the language. In this paper, we will demonstrate that, even after disassociated from their bindings to particular words, the “shape” of these point clouds reflect the history of the languages they represent, by using techniques from topological data analysis (TDA), a field studying spatial aspects of data.
The history of language evolution is typically studied by identifying cognate word pairs of languages $L_{1}$ and $L_{2}$, namely homologous word forms inherited from the common ancestor of these languages. Cognates typically share some level of surface-form similarity between them, with language pairs sharing more cognates presumed to be more closely related. The identification of cognate pairs traditionally requires extensive labor by skilled linguists, although computational methods for detecting cognate pairs are emerging Kondrak (2001); Lefever et al. (2020); St Arnaud et al. (2017).
An interesting aspect of the topological analysis of word embeddings we consider here is that it makes no assumption or use of translation pairs, surface similaries between words, or cognate pairs. Word embeddings are treated as unlabeled sets of points, shorn of any binding to corresponding tokens. TDA compares the relative shape of word embeddings in terms of structural properties like components and holes. That we will demonstrate even a weak ability to reconstruct language phylogenies through TDA is surprising, and suggests that there are hitherto unknown structures reflected in word embeddings that go beyond the semantics of individual words.
@@
  • Comparing the Shapes of Word Embeddings using Topological Data Analysis (TDA) – Through our experimental setup on language phylogeny reconstruction, we will test two related properties concerning the geometric structure of word embeddings. First, we will show that the shape of the unlabeled word embedding of a language carries information about its history and structure. Second, we show that this information can be at least partially recovered through persistent homology, a standard tool in TDA.
  • Statistical Evaluation of TDA-Based Language Phylogenies – Evaluating the information content of language distance matrices inferred using TDA is best done by converting these matrices into phylogenic trees, and then assessing the quality of these trees against a reference standard. Representative trees constructed using our methods are presented in Figure 2. We perform an extensive study evaluating the accuracy of TDA distance matrices derived from two different types and three different dimensions of persistence diagram, compared using four different distances between persistence diagrams. Each of these 24 different TDA matrices was evaluated on two popular tree construction algorithms (UPGMA and neighbor joining) and compared against a gold-standard language tree (from Ethnologue) under six different tree-similarity metrics.
    Certain matrices and metrics prove less capable of encoding or recognizing the structural similarities inherent in word embeddings. But permutation tests show that for 484 out of 864 combinations of parameters, the quality of the reconstructed trees were significant to the 0.05-level, and many substantially stronger than that. Even stronger bounds come in the number of standard deviations our TDA trees are closer to the gold-standard tree compared to the mean statistical background. Many of our TDA-based trees sit from four to seven $\sigma$ from the mean. Under assumptions of normality, the maximum achieved $6.87\sigma$ corresponds to a Bonferroni-corrected $p$-value of $2.77\times 10^{-9}$ of seeing a result like this by chance.
@@
  • A Case Study in Topological Data Analysis – TDA provides a relatively new set of tools for data science able to capture features often missed by more traditional approaches. Due to the general framework it provides, there are many parameters to choose from—selecting the filtration function on data, dimension of studied features, distance notion between the extracted features, etc.—and a priori no obvious canonical choices. Our analysis on word embeddings represents a rigorous evaluation for an interesting use-case of this technology. We demonstrate that TDA methods successfully capture aspects of the shape of language embeddings. Our analysis shows that persistence diagrams in dimensions 2, which are not that often used in large high-dimensional data because of the expense in constructing them and difficulty to interpret them, yield statistically-strong results in our experiments, in some cases outperforming dimensions 0 and 1.
@@
We presume a reader from computational linguistics, with a basic familiarity with the concepts and literature of word embeddings and language phylogenies, but no prior exposure to topological data analysis. Our paper is organized as follows. Important TDA concepts like persistence diagrams and the Victoris-Rips complex are reviewed in Section 2. Algorithms for constructing language phylogenies from distance matrices and evaluating the resulting trees are presented in Section 3. We describe our TDA-based analysis pipeline in Section 4 with computational results reported in Section 5.
@@
Our primary goal in this work is to initiate the study of the shape of languages via word embeddings, not to advance the state-of-the-art in language phylogeny reconstruction. Indeed, cognate-based analysis should clearly dominate our unlabeled point methods except perhaps in pathological situations, such as completely undeciphered written languages. But we do believe that our new TDA-based tools hold promise for richer computational language analysis, and raise interesting question about which properties in language-space the topological features we employ correspond to.
@@ 2 Topological Data Analysis
Topological data analysis (TDA) is a growing field that applies methods developed for studying shapes to data, both geometrical and abstract, to extract features that are often not captured by classical approaches of data science. In the case of points or vectors in $2$- or $3$-dimensional space, we observe certain patterns like clusters of points, empty spaces surrounded by points (voids) or tunnels in 3D. TDA studies such structures in a quantitative way.
To keep the explanation intuitive, we stay with low dimensional point sets, but the general framework is limited neither by dimension of the data nor by the Euclidean metric—indeed, we use $300$-dimensional word embedddings and impose both Euclidean and cosine metrics. For a gentle introduction to TDA see Edelsbrunner (2014). For more technical details, see Edelsbrunner and Harer (2010). For applications of TDA to a wide variety of real-world data, see the database of non-theoretical uses of topology Giunti et al. (2022).
@@ 2.1 Persistent homology
We first describe the idea of persistent homology in a simple intuitive setting—consider a data set of points in a Euclidean plane. To define the ‘shape’ of such a set, choose a radius, place a disk of that radius centered at each point, and consider the union of the disks; see Figure 1. Two simple descriptors for such a shape are the number of connected components and the number of holes. These are Betti numbers111 The formal definition of Betti numbers depends on the choice of coefficients for homology. We use the field of size two, $\mathbb{Z}/2\mathbb{Z}$, which is the standard choice for persistent homology. of the shape, $\beta_{0}$ and $\beta_{1}$, respectively—a notion formally defined by homology, a tool in the mathematical field of algebraic topology; see, e.g., Hatcher (2002).
@@
For different radii of the disks, we get different Betti numbers. Consider growing the radius from zero to infinity. We can plot the Betti numbers depending on the radius, and treat the curves, $\beta_{0}(r)$, $\beta_{1}(r)$, as descriptors for the shape of our point set. The problem is that such a curve can differ a lot when we slightly perturb the points. For a stable descriptor, we turn to persistent homology, which, in addition, pairs the birth and death radii of each component or hole. For example, there is a minimum radius $r_{b}$ for which the disks enclose a space in the plane, separating it from its surroundings. This is the birth radius of a hole. Then there is a larger minimum radius $r_{d}$, for which the area we previously enclosed is fully covered by the disks. The interval $\left[r_{b},r_{d}\right)$, often called a bar, is a single feature recorded by persistent homology. It is the interval of radii during which this particular hole contributes “plus one” to the Betti number $\beta_{1}$.
@@
We collect all such intervals. Each interval is described by its two endpoints, so we can plot them all in a scatter plot with axes “birth radius” and “death radius”; see Figure 1(c). This plot is the persistence diagram of the data point set. For a point set in a plane, we get two persistence diagrams: one for components of connectivity, i.e., the 00-dimensional features contributing to $\beta_{0}$, and one for ‘holes’ or ‘loops’, i.e., the $1$-dimensional features contributing to $\beta_{1}$. Usually we plot both in the same figure, separating them by color and shape of the markers.
Refer to captionRefer to captionRefer to caption
@@
Figure 1: Unions of disks centered at data points in a Euclidean plane define a ‘shape’ that varies with the radius of the disks. Persistent homology studies different types of ‘gaps’ within this growing shape, e.g., 00-dimensional gaps between components of connectivity, and $1$-dimensional holes within the shape. Each such feature is born at some radius, and dies at some later radius, spanning an interval $[r_{b},r_{d})$. Those intervals are summarised in a persistence diagram—each interval is represented by a dot with coordinates $(r_{b},r_{d})$. The left and middle figures show the growing disks around a point set $A\subseteq\mathbb{R}^{2}$ with the radii of the endpoints of an interval $[r_{b},r_{d})=[0.072,0.119)$, as well as the Vietoris-Rips complexes $\mathrm{VR}_{r_{b}}({A})$, $\mathrm{VR}_{r_{d}}({A})$. On the right is the persistence diagram of the growing disks with highlighted blocks corresponding to radii $r_{b}$, $r_{d}$.
@@
When we increase the dimension of the data, we can have more different kinds of features. A $2$-dimensional feature is an uncovered space completely surrounded by covered space—like a hole completely inside a Swiss cheese. Analogously to loops in 3D, in higher dimensional data, a $2$-dimensional feature would not be a completely enclosed space, but rather a drawing of a $2$-dimensional sphere that cannot be contracted into a point. Even though our actual data is $300$-dimensional, we only consider features of dimension 00, $1$ and $2$, mainly for computational reasons. Examples of diagrams used in our analysis are provided in Appendix Figure 9.
Persistent homology is stable. That is, if the data points are perturbed a bit, the points on the persistence diagrams are also only perturbed a bit Cohen-Steiner et al. (2007). Note that points can appear and disappear on the diagonal of a persistence diagram in the process. To properly state the stability, we need to define a notion of distance that reflects that. Usually the stability is phrased with respect to the bottleneck distance or $1$-Wasserstein distance—see Section 2.3 for the definitions.
2.2 Vietoris-Rips complex
@@ 2.3 Distances between persistence diagrams
In this section, we describe several established notions of distances between persistent diagrams. The list we use is by no means exhaustive—see, e.g., Ali et al. (2023) for a survey of methods, and Carrière et al. (2020) for an approach unifying many of those methods under a common framework. We chose four distances that are widely used in theory or in applications and have readily available implementations. Two of those are defined directly on the diagrams, while the other two first vectorize each diagram and then compare the vectors with the standard Euclidean metric.
We fix two persistent diagrams, $\mathrm{PD}_{1}$ and $\mathrm{PD}_{2}$, describing topological features of the same dimension. An element $r$ of $\mathrm{PD}_{i}$ is a point in a plane representing an interval, and we will denote its endpoints by $r=[r_{b},r_{d})$ for “birth” and “death”.
@@
Bottleneck distance is a classical notion, but it is computationally costly since we have to search the space of matchings.
@@
This distance takes each interval in either diagram into consideration, not just the worst match as the bottleneck distance. It is stable with respect to data perturbations Cohen-Steiner et al. (2010); Skraba and Turner (2023). In our analysis, we use a computationally more viable variant—sliced Wasserstein distance which approximates $1$-Wasserstein distance Carrière et al. (2017).
@@
Rather than directly comparing the diagrams, we can first transform each into a vector and then compare the vectors. One approach to vectorize a diagram is to blur the points and then treat the plot as a raster image, called persistence image Adams et al. (2017). Persistence images are stable and approximate the $1$-Wasserstein distance.
Persistence images are easy to compute, but there is several parameters we need to fix—the size of the pixels and the grid, the $\sigma$-parameter of the Gaussians, and the weight function controlling the heights of the Gaussians based on the $d-b$ parameter.
@@
3 Constructing/Evaluating Phylogenies
We emphasize that reconstructing language phylogenies is an enormous area within computational linguistics, and space does not permit us to make a substantial survey of work in this field. See Dunn (2015); Pereltsvaig and Lewis (2015); Pompei et al. (2011) for recent reviews of this literature. We are interested in phylogeny reconstruction primarily as a task to evaluate the extent to which the unlabeled topological “shape” of word embeddings retain and reflect broader properties of the underlying languages. Thus we use standard distance-based tree reconstruction algorithms (unweighted pair group method with arithmetic means (UPGMA) and neighbor joining) Gusfield (1997) and tree similarity metrics detailed below.
3.1 Data
The historical origin of languages has been extensively studied, and much is known, but debate remains vigorous even for the Indo-European languages we study here Gray and Atkinson (2003); Pereltsvaig and Lewis (2015); Longobardi et al. (2013). The reference tree we use comes from Ethnologue (https://www.ethnologue.com), which is the most widely consulted inventory of the world’s languages Lewis et al. (2024). First published in 1951 and now in its 26th edition, Ethnologue currently records data about 7,168 living languages. Although there is no universal agreement on language origins (see e.g. (Hammarström, 2015) for a critique of Ethnologue), it provides a broadly acceptable reference tree for our evaluation purposes, particularly for the major Indo-European languages we consider here.
FastText Bojanowski et al. (2017) (https://fasttext.cc/) provides a set of pre-trained, 300-dimensional word embeddings for 157 languages, trained on Common Crawl and Wikipedia. There are 81 Indo-European languages that appear in the Ethnologue reference tree and also have pre-computed FastText embeddings. To protect against the risk that low-resource languages might have less reliable embeddings, we will also evaluate filtrations to the most popular 30 and 50 languages—see Section 4. The Ethnologue reference trees for all filtrations are provided for inspection in Appendix Figure 7.
3.2 Constructing Algorithmic Phylogenies
In this paper, we experiment with two popular agglomerative hierarchical clustering algorithms for reconstructing phylogenetic trees from a pairwise distance matrix, $d(\cdot,\cdot)$. We start with each language as an individual cluster, and then we connect two “closest” clusters in each step. The difference is what “closest” means for each of the algorithms:

### v2

# The Shape of Word Embeddings: Quantifying Non-Isometry With Topological Data Analysis

@@ Abstract
Word embeddings represent language vocabularies as clouds of $d$-dimensional points. We investigate how information is conveyed by the general shape of these clouds, instead of representing the semantic meaning of each token. Specifically, we use the notion of persistent homology from topological data analysis (TDA) to measure the distances between language pairs from the shape of their unlabeled embeddings. These distances quantify the degree of non-isometry of the embeddings. To distinguish whether these differences are random training errors or capture real information about the languages, we use the computed distance matrices to construct language phylogenetic trees over 81 Indo-European languages. Careful evaluation shows that our reconstructed trees exhibit strong and statistically-significant similarities to the reference.
The Shape of Word Embeddings: Quantifying Non-Isometry With Topological Data Analysis
Ondřej Draganov Institute of Science and Technology Austria (ISTA) Am Campus 1, 3400 Klosterneuburg, Austria ondrej.draganov@ist.ac.at                        Steven Skiena Dept. of Computer Science Stony Brook University Stony Brook, NY 11794-2424 USA skiena@cs.stonybrook.edu
@@ 1 Introduction
Does the shape of an unlabeled, monolingual word embedding carry relevant information about the language it represents? Word embeddings are well-established objects of interest in natural language processing, being $d$-dimensional vector representations that capture the semantics of each vocabulary word. The vocabulary of language $L$ can thus be viewed as a cloud of points, whose geometric and structural properties encode considerable information about the language. In this paper, we demonstrate that, even after disassociated from their bindings to particular words, the “shape” of these point clouds correlates with the history of the languages they represent. We use techniques from topological data analysis (TDA), a field studying spatial aspects of data, to quantify this correlation.
How much shapes of word embeddings differ, and for what reasons, is an ongoing debate with important consequences. Many authors assume that the word embeddings of different languages are essentially isometric: meaning they only differ by an orthogonal transformation. Such assumptions are necessary, for example, to construct translators between languages via bilingual lexicon induction (Miceli Barone, 2016; Lample et al., 2018; Zhang et al., 2017). This assumption has been disputed by Søgaard et al. (2018) or Patra et al. (2019). Later Vulić et al. (2020) argued that observed deviations from isometry are largely due to insufficient data or training in construction of embeddings for some languages. In this paper, we demonstrate that these observed non-isometries across languages contain real information and are thus not random training artifacts. We analyze FastText embeddings of 81 Indo-European languages and quantify non-isometry among the 10k most frequent tokens of each language with TDA-based distance matrices. To show that those distances contain real information we construct phylogenetic trees and compare these to a gold-standard reference tree from Ethnologue Lewis et al. (2024).
The history of language evolution is typically studied by identifying cognate word pairs of languages $L_{1}$ and $L_{2}$, namely homologous word forms inherited from the common ancestor of these languages. Cognates usually share some level of surface-form similarity between them, with language pairs sharing more cognates presumed to be more closely related. The identification of cognate pairs traditionally requires extensive labor by skilled linguists, although computational methods for detecting cognate pairs are emerging Kondrak (2001); Lefever et al. (2020); St Arnaud et al. (2017).
An interesting aspect of TDA on word embeddings is that it makes no assumption or use of translation pairs, surface similarities between words, or cognate pairs. Word embeddings are treated as unlabeled sets of points, shorn of any binding to corresponding tokens. TDA compares the relative shape of word embeddings in terms of structural properties like components of connectivity and holes. That we demonstrate even a weak ability to reconstruct language phylogenies through TDA is quite surprising, and suggests that there are hitherto unknown structures reflected in word embeddings that go beyond the semantics of individual words.
@@
Further evidence that even within the Indo-European family the non-isometry between unlabeled word embeddings is not completely erroneous.
The fact that we reconstruct phylogenetic trees that, despite not very precise, are considerably better than chance, suggests that there is some real information represented in the differences between unlabeled shapes of embeddings. It is fair to argue that more careful training, normalization and lemmatisation could yield more similar embeddings that miss the information; were that the case, it would suggest that the differences mitigated by those processes correlate with language history, which is an equally interesting observation.
A case study of natural generalizations of previously used methods to quantify non-isometry.
@@
TDA provides a relatively new set of tools for data science able to capture features often missed by more traditional approaches. Due to the general framework it provides, there are many parameters to choose from and a priori no obvious canonical choices. Our analysis on word embeddings represents a rigorous evaluation for an interesting use-case of this technology.
A particular instance of TDA was used in Patra et al. (2019); Vulić et al. (2020) as an approximation Chazal et al. (2009) of difficult-to-compute Gromov-Hausdorff distance to quantify non-isometry of word embeddings. In the TDA terminology, they computed bottleneck distance between degree 00 persistent diagrams—we additionally explore degrees $1$ and $2$, and three different distances in addition to bottleneck; each combination for both Euclidean and cosine metric on the word embedding. Altogether this totals to 24 different combinations of parameters, and for each we obtain a distance matrix between the studied languages. All those notions of distances are invariant under orthogonal transformations [^1] Cosine distances can change with translation—a shift of the whole point cloud., and therefore serve as quantifications for non-isometry. The alternative combinations can be better suited for studying word embeddings, e.g., by being less prone to irrelevant outliers that can artificially significantly increase the GH distance. Our analysis also shows that degree $2$, which is not that often used in large high-dimensional data because of the expense in constructing them and difficulty to interpret them, yield statistically-strong results in our experiments, in some cases outperforming degrees 0 and 1.
Statistical Evaluation of TDA-Based Language Distances with Phylogenetic Trees.
Evaluating the information content of language distance matrices inferred using TDA is best done by converting these matrices into phylogenetic trees, and then assessing the quality of these trees against a reference standard. Each of 24 different TDA matrices, constructed based on different combinations of parameters, was evaluated on two popular tree construction algorithms—UPGMA and neighbor joining—and compared against a gold-standard language tree from Ethnologue under six different tree-similarity metrics.
Certain choices of parameters prove less capable of recognizing the structural similarities inherent in word embeddings. But permutation tests show that for 484 out of 864 combinations of parameters [^2] See Section 4 for a break-down of the $864=2\cdot 3\cdot 4\cdot 3\cdot 2\cdot 6$ combinations of parameters., the quality of the reconstructed trees were significant to the 0.05-level, and many substantially stronger than that. Even stronger bounds come in the number of standard deviations our TDA trees are closer to the gold-standard tree compared to the mean statistical background. Many of our TDA-based trees sit from four to seven $\sigma$ from the mean. Under assumptions of normality, the maximum achieved $6.87\sigma$ corresponds to a Bonferroni-corrected $p$-value of $2.77\times 10^{-9}$ of seeing a result like this by chance.
@@
We presume a reader from computational linguistics, with a basic familiarity with the concepts and literature of word embeddings and language phylogenies, but no prior exposure to topological data analysis. Our paper is organized as follows. Important TDA concepts like persistence diagrams are reviewed in Section 2. Algorithms for constructing language phylogenies from distance matrices and evaluating the resulting trees are presented in Section 3. We describe our TDA-based analysis pipeline in Section 4 with computational results reported in Section 5, and conclude with Section 6.
@@
Our primary goal in this work is to initiate the study of the shape of languages via word embeddings, not to advance the state-of-the-art in language phylogeny reconstruction. Indeed, cognate-based analysis should clearly dominate our unlabeled point methods except perhaps in pathological situations, such as completely undeciphered written languages. But we do believe that the new TDA-based tools hold promise for richer computational language analysis, and raise interesting questions about which properties in language-space the topological features we employ correspond to.
@@ 2 Topological Data Analysis
Topological data analysis (TDA) is a growing field that applies methods developed for studying shapes to data, both geometrical and abstract, to extract features that are often not captured by classical approaches of data science. In the case of points or vectors in $2$D or $3$D, we observe patterns like clusters of points, loops, voids or tunnels.
In this paper, the input is a word embedding of a language—collection of points in high-dimensional Euclidean space, one for each token—and the output is a summary of its topological features. We compute such summaries for embeddings of different languages, and use them to compare the structural similarity. As orthogonal transformations would preserve the summaries, those differences quantify non-ismoetry of the different embeddings.
To keep the explanation intuitive, we stay with low dimensional point sets, but the general framework is limited neither by dimension of the data nor by the Euclidean metric—indeed, we use $300$-dimensional word embedddings and impose both Euclidean and cosine metrics. For a gentle introduction to TDA see Edelsbrunner (2014). For more technical details, see Edelsbrunner and Harer (2010). For applications of TDA to a wide variety of real-world data, see the database of non-theoretical uses of topology (Giunti et al., 2022).
@@ 2.1 Persistent Homology
We describe the idea of persistent homology in a simple intuitive setting—a data set of points in a Euclidean plane. To define the ‘shape’ of such a set, choose a radius, place a disk of that radius centered at each point, and consider the union of the disks; see Figure 1. Two simple descriptors for such a shape are the number of connected components and the number of holes—Betti numbers333 The formal definition of Betti numbers depends on the choice of coefficients for homology. We use the field of size two, which is the standard choice for persistent homology. $\beta_{0}$ and $\beta_{1}$, respectively. The notion is formally defined by homology from the mathematical field of algebraic topology; see, e.g., Hatcher (2002).
@@
For different radii of the disks, we get different Betti numbers. Consider growing the radius from zero to infinity. We can plot the Betti numbers depending on the radius, and treat the curves, $\beta_{0}(r)$, $\beta_{1}(r)$, as descriptors for the shape of our point set. The problem is that such a curve can differ a lot when we slightly perturb the points. For a stable descriptor, we turn to persistent homology, which, in addition, pairs the birth and death radii of each component or hole. Components are all born at radius 0 and each time two components merge at radius $r$, we get a birth-death pair $[0,r)$. For a hole or a loop, there is a minimum birth radius, $r_{b}$, for which the disks enclose a space in the plane, separating it from its surroundings, and there is a larger minimum death radius, $r_{d}$, for which the area we previously enclosed is fully covered by the disks. The interval $\left[r_{b},r_{d}\right)$, often called a bar, is a single feature recorded by persistent homology. It is the interval of radii during which this particular hole contributes “plus one” to the corresponding Betti number.
@@
We collect all such intervals. Each interval is described by its two endpoints, so we can plot them all in a scatter plot with axes “birth radius” and “death radius”; see Figure 1(c). This plot is the persistence diagram of the data. For a point set in a plane, we get two persistence diagrams: one for components of connectivity—degree 0—and one for ‘holes’ or ‘loops’—degree 1. Usually we plot both in the same figure, separating them by color or shape of the markers.
@@
Figure 1: Unions of disks centered at data points in a Euclidean plane define a ‘shape’ that varies with the radius of the disks. In our setting, each point is one token of a word embedding, and the space is $\mathbb{R}^{300}$ rather than $\mathbb{R}^{2}$. Persistent homology studies different types of ‘gaps’ within this growing shape—00-dimensional gaps between components of connectivity, and $1$-dimensional holes within the shape. Each such feature is born at some radius, and dies at some later radius, spanning an interval $[r_{b},r_{d})$. Those intervals are summarised in a persistence diagram; each interval is represented by a dot with coordinates $(r_{b},r_{d})$. The left and middle figures correspond to birth and death radii of the loop in the lower-right corner: $[r_{b},r_{d})=[0.072,0.119)$. Overlaid is the Vietoris-Rips complex discretizing the pink shape to allow for computations (see A.1). On the right is the persistence diagram of the growing disks with highlighted blocks corresponding to radii $r_{b}$, $r_{d}$.
@@
When we increase the dimension of the data, we can have more different kinds of features. In 3D, we grow balls rather than disks around the points. A degree 2 feature is an uncovered space completely surrounded by covered space—like a hole completely inside a Swiss cheese. A degree 1 feature is a loop drawn into the union of the balls that cannot be contracted to a point within that union. As loops in 3D, a $2$-dimensional feature in higher dimensional data would not be a completely enclosed space, but rather a drawing of a $2$-dimensional sphere that cannot be contracted to a point.
Although our actual data is $300$-dimensional, we only consider features of degree 00, $1$ and $2$, mainly for computational reasons. Examples of diagrams from our analysis are in Appendix Figure 9.
Persistent homology is stable against perturbations. That is, if the data points are perturbed a bit, the points on the persistence diagrams are also only perturbed a bit Cohen-Steiner et al. (2007). Note that in contrast, it is not very stable against outliers, which is a limitation of the tool that can be somewhat mitigated with the right choice of distance between persistence diagrams.
@@ 2.2 Distances Between Persistence Diagrams
In this section, we describe several established notions of distances between persistent diagrams. The list we use is by no means exhaustive—see, e.g., Ali et al. (2023) for a survey of methods, and Carrière et al. (2020) for an approach unifying many of those methods under a common framework. We chose four distances that are widely used in theory or in applications and have readily available implementations. Two of those are defined directly on the diagrams, while the other two first vectorize each diagram and then compare the vectors with the standard Euclidean metric. Note two properties that make defining distances somewhat involved: firstly, different persistence diagrams have generally different number of points with no canonical matching between them; secondly, slightly perturbing data can create or remove points at the diagonal, which needs to be taken into consideration.
We fix two persistent diagrams, $\mathrm{PD}_{1}$ and $\mathrm{PD}_{2}$, describing topological features of the same degree. An element $r$ of $\mathrm{PD}_{i}$ is a point in a plane representing an interval, and we will denote its endpoints by $r=[r_{b},r_{d})$ for “birth” and “death”.
@@
Bottleneck distance is a classical notion, but it is computationally costly since we have to search the space of matchings. It is comparing the similarity of the diagrams “on the nose”. This is well-suited for those applications in which objects are considered similar only if their “hole-structure” is almost identical, but can be too penalising when we are interested more vaguely in the distribution of different kinds of holes. It is also very sensitive to outliers.
@@
This distance takes each interval in either diagram into consideration, not just the worst match as the bottleneck distance. For discussion on stability with respect to data perturbations see Cohen-Steiner et al. (2010); Skraba and Turner (2023). In our analysis, we use a computationally more viable variant—sliced Wasserstein distance which approximates $1$-Wasserstein distance Carrière et al. (2017).
@@
Rather than directly comparing the diagrams, we can first transform each into a vector and then compare the vectors. One approach to vectorize a diagram is to blur the points and then treat the plot as a raster image, called persistence image Adams et al. (2017). Persistence images are stable and approximate the $1$-Wasserstein distance. More formally, we put a two-dimensional Gaussian over each dot in the persistence diagram, and consider their sum. Then we overlay a grid over the domain, and compute the integral of the sum of Gaussians over each grid square. This yields a value for a pixel. Since the dots further away from the diagonal are more relevant, there is a weight function defining the height of the Gaussian based on the distance away from the diagonal. The distance between two diagrams is then the Euclidean distance between their persistence images viewed as vectors.
Persistence images are easy to compute, but there is several parameters we need to fix—the size of the pixels and the grid, the $\sigma$-parameter of the Gaussians, and the weight function controlling the heights of the Gaussians based on the distance from the diagonal.
@@
3 Constructing Phylogenetic Trees
To evaluate whether the distances between languages calculated based on persistence homology are relevant, we use this data to attempt to reconstruct the historical evolution of Indo-European languages. We perform a hierarchical clustering, obtain a phylogenetic tree, and compare this tree to a ground-truth reference tree.
We emphasize that reconstructing language phylogenies is not the goal of our analysis, but rather we use it as a tool to evaluate whether the computed distances carry a relevant signal, meaning the unlabeled topological “shape” of word embeddings retain and reflect broader properties of the underlying languages. Phylogeny reconstruction is an enormous area within computational linguistics, and space does not permit us to make a substantial survey of work in this field. See Dunn (2015); Pereltsvaig and Lewis (2015); Pompei et al. (2011) for recent reviews of this literature.
In this paper, we experiment with two popular agglomerative hierarchical clustering algorithms for reconstructing phylogenetic trees from a pairwise distance matrix, $d(\cdot,\cdot)$. We start with each language as an individual cluster, and then we connect two closest clusters in each step. The difference is what “closest” means for each of the algorithms: